five years ago the combined age of six girls was 67. what will be their combined age in four yearss time?
Found 4 solutions by Alan3354, josgarithmetic, ikleyn, TeachMath:
Answer by Alan3354(69443)

Think about it.

Answer by josgarithmetic(39620)
Answer by ikleyn(52813)

.
If 5 years ago the combined age of 6 girls was 67, then their combined present age is 67 + 6*5 = 97 years.


Then in 4 years, their combined age will be 97 + 4*6 years.


Answer by TeachMath(96)


67 + 9(6) = 67 + 54 = 121 yaars